OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x – 3y – 3z = –13 x + 2y – z = –1 2x + 3y + 2z = 0

OpenStudy (anonymous):

x – 3y – 3z = –13 (x + 2y – z = –1)*3 *************** x – 3y – 3z = –13 3x + 6y – 3z = –3 ----------------------subtract (x-3x) + (-3y-6y) + (-3z - -3z) = (-13- -3) -2x -9y + 0z = -10 -2x - 9y = -10 now you try the second and third equations

OpenStudy (anonymous):

I'm so lost right now!

OpenStudy (anonymous):

why, what dont you understand?

OpenStudy (anonymous):

hmmmmmmm................EVERYTHING that has to do with math!

OpenStudy (anonymous):

lol why? math is easy and sweet lol

OpenStudy (anonymous):

worst subject for me

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

11th

OpenStudy (anonymous):

you will have to multiply one or two of the equations with numbers so when you add or subtract them one of the terms drop in equation 1 and 2 you needed to drop z 1st eq had -3z 2nd eq had -z in order to drop that term we need to multiply the 2nd eq by 3